### v1

@@ 4.2.2 Strongly reticulation-visible networks
We quickly summarize the collection of results obtained so far. By Lemma 12, bridge-nodes are stable. Therefore, if $v_{b}$ is a bridge-node, then by Corollary 2, $\mu(v_{b})\leq 2$. Furthermore, by Lemma 8, $\mu(v_{b})$ does not belong to a pair of tree-clones. Finally, all leaves are bridge-nodes, therefore all unit vectors in $\mu(\mathcal{N})$ belong to bridge-nodes, so we only still have to consider $\mu$-vectors which are not unit vectors. Now let $\mathcal{N}=(V,E)$ be a semi-binary stack-free network such that all tree-nodes in $V$ have at most 2 children. We use the following lemmas (Lemma 19 to Lemma 24) to show how to determine whether a non-unit $\mu$-vector belongs to a bridge-node (Proposition 3).
@@
Given a tree-node $v_{b}$, let us first assume $v_{b}$ has two children $c_{1}$ and $c_{2}$ which are reticulations. Because $c_{1}$ and $c_{2}$ are reticulations they must have at least one other parent besides $v_{b}$. Neither $c_{1}$ is the parent of $c_{2}$ nor is $c_{2}$ the parent of $c_{1}$ because the network is stack-free. Note that, not both $c_{1}$ and $c_{2}$ can have a parent which is a descendant of the other one, because in that case there would be a cycle, from $c_{1}$ to the parent of $c_{2}$ to $c_{2}$ to the parent of $c_{1}$ to $c_{1}$. Furthermore, all descendants of $v_{b}$ except $v_{b}$ itself are descendants of $c_{1}$ or $c_{2}$. Therefore, either $c_{1}$ or $c_{2}$ has a parent which is not a descendant of $v_{b}$. But then there would be a path from the root to a leaf below $v_{b}$ via this parent, which does not pass through $v_{b}$. Thus, by Lemma 12, this means $v_{b}$ is not a bridge-node.
@@
We will show the lemma is true with a proof by contradiction, by showing that there cannot be a second pair. Assume $\mu(v_{b})$ is a bridge-node which is not a leaf and $\mu(\mathcal{N})$ contains at least two pairs $\mu(x),\mu(y)$ and $\mu(k),\mu(\ell)$ such that $\mu(v_{b})=\mu(x)+\mu(y)=\mu(k)+\mu(\ell)$. W.l.o.g. we can assume $\mu(x)$ and $\mu(y)$ belong to the children $x,y$ of $v_{b}$. By Lemma 19, we know that $\mu(k)\neq\mu(x)$, because otherwise $\mu(\ell)=\mu(y)$, in which case both #$\#\mu(x)\geq 2$ and #$\#\mu(y)\geq 2$, and if $\mu(x)=\mu(y)$ then #$\#\mu(x)\geq 4$. The same goes for $\mu(k)\neq\mu(y)$, $\mu(\ell)\neq\mu(x)$ and $\mu(\ell)\neq\mu(y)$. Note also that none of $x,y,k,\ell$ are ancestors of $v_{b}$ because their $\mu$-vectors are each lower than $\mu(v_{b})$. Now note that, by Corollary 3, $\mu(v_{b})>\mu(k)$ and $\mu(v_{b})>\mu(\ell)$ implies they are descendants of $v_{b}$. However, as we mentioned $x,y$ are the children of $v_{b}$. This means $k,\ell$ must be descendants of $x,y$, which are not equal to $x,y$, because their $\mu$-vectors differ. From this it follows that $\mu(x)+\mu(y)>\mu(k)+\mu(\ell)$. This contradicts our assumption that the sums were equal. ∎
@@
If $\mu(v_{t})\in\mu(\mathcal{N})$ belongs to a tree-clone, then there are at least two tree-nodes $v_{1}$ and $v_{2}$ with $\mu$-vector equal to $\mu(v_{1})$. By Observation 1, unit-vectors do not belong to tree-clones, therefore $v_{1}$ and $v_{2}$ are not leaves. This means $v_{1}$ and $v_{2}$ each have two children. Let $c_{1}$ and $c_{2}$ be the children of $v_{1}$, and let $c_{3}$ and $c_{4}$ be the children of $v_{2}$. Note that $c_{1}$ must be distinct from $c_{2}$, and $c_{3}$ must be distinct from $c_{4}$, because phylogenetic networks do not contain parallel edges. We then have: $\mu(c_{1})+\mu(c_{2})=\mu(v_{1})=\mu(v_{2})=\mu(c_{3})+\mu(c_{4})$. The statement can be broken down into three unique cases.
@@
Let $\mu(v_{b})\in\mu(\mathcal{N})$ belong to a bridge-node $v_{b}$. Let $I_{b}$ be the set of leaves below $v_{b}$, so that $\mu_{i}(v_{b})=0$ for $i\notin I_{b}$. For any $\mu(x)\in\mu(\mathcal{N})$ with $\mu(x)\not<\mu(v_{b})$, it holds that $\mu(x)=P_{xv_{b}}\mu(v_{b})+\mu^{\prime}(x)$, where $P_{xv_{b}}$ is a non-negative integer, equal to the number of paths from $x$ to $v_{b}$, and $\mu^{\prime}(x)$ is a $\mu$-vector such that $\mu^{\prime}_{i}(x)=0$ for $i\in I_{b}$.
@@
Let $\mu(v_{b})\in\mu(\mathcal{N})$ belong to a bridge-node. For any $\mu(x)\in\mu(\mathcal{N})$ with $\mu(x)\not<\mu(v_{b})$, by Lemma 12, $\mu(x)$ does not belong to a descendant of the bridge-node with $\mu$-vector $\mu(v_{b})$. Therefore, all paths from $x$ to leaves below $v_{b}$ must visit $v_{b}$. Thus, if $a$ is a leaf below $v_{b}$, then any path from $x$ to $a$ is the composition of a path from $x$ to $v_{b}$ and a path from $v_{b}$ to $a$. It follows that for each such leaf $a$, $\mu_{a}(x)=P_{xv_{b}}\mu_{a}(v_{b})$. From this it follows that $\mu(x)=P_{xv_{b}}\mu(v_{b})+\mu^{\prime}(x)$, where $\mu^{\prime}(x)$ contains only the paths to leaves not below $v_{b}$, and therefore $\mu^{\prime}(x)$ is a $\mu$-vector such that $\mu^{\prime}_{i}(x)=0$ for $i\in I_{b}$. ∎
@@
We will show this with a proof by contradiction. Let us assume that $\mu(v_{b})\in\mu(\mathcal{N})$ belong to a bridge-node $v_{b}$ and $\mu(\mathcal{N})$ contains some $\mu(x),\mu(y)$ such that $\mu(x)<\mu(v_{b})$. Now let us assume, $\mu(\mathcal{N})$ contains $\mu(k),\mu(\ell)$ with $\mu(k)\not<\mu(v_{b})$ and $\mu(\ell)\not<\mu(v_{b})$, such that $\mu(k)+\mu(\ell)=\mu(x)+\mu(y)$. First note, that if $\mu(y)\leq\mu(v_{b})$, then $\mu(y)$ belongs to a descendant of $v_{b}$ and therefore $\mu(k)+\mu(\ell)=\mu(x)+\mu(y)\leq\mu(v_{b})$. But this contradicts our assumption that $\mu(k)\not<\mu(v_{b})$. Therefore, we can assume that $\mu(y)\nleq\mu(v_{b})$. Then, by Lemma 22:
@@
We will argue by contradiction. Assume $\mu(\mathcal{N})$ contains $\mu(v_{b}),\mu(z),\mu(x)$ and $\mu(y)$ as described in the lemma and let $\mu(v_{b})$ belong to a bridge-node $v_{b}$. If $\mu(z)$ belongs to a reticulation $r$, then the child of $r$ must be a tree-node with the same $\mu$-vector, so w.l.o.g. we can assume $\mu(z)$ belongs to a tree-node $z$. Because $\mu(z)=\mu(x)+\mu(y)$, we know $z$ is not a leaf, because $\mu(z)$ is not a unit vector. If $\mu(x),\mu(y)$ is the only pair in $\mu(\mathcal{N})$ which sum up to $\mu(z)$ then $z$ must be the parent of $x$. If there are more pairs in $\mu(\mathcal{N})$ then by Lemma 23, each of those pairs must contain at least one $\mu$-vector lower than $\mu(v_{b})$. This means that in any case $z$ will have a child with $\mu$-vector lower than $\mu(v_{b})$. However, $\mu(z)\nleq\mu(v_{b})$ implies that $z$ is not a descendant of $v_{b}$. Then there must be a path from the root to a leaf below $v_{b}$ via $z$ which does not visit $v_{b}$. By Lemma 12, this means that $v_{b}$ is not a bridge-node. Which contradicts our assumption. ∎
Observation 2.
Proposition 3.
@@
  • $\mu(k)$ and $\mu(\ell)$ do not both have multiplicity greater than 1 in $\mu(\mathcal{N})$ unless $\mu(k)=\mu(\ell)$ with #$\#\mu(k)=2$, and

### v2

@@ 4.2.2 Strongly reticulation-visible networks
We quickly summarize the collection of results obtained so far. By 2, bridge-nodes are stable. Therefore, if $v_{b}$ is a bridge-node, then by 2, #$\#\mu(v_{b})\leq 2$. Furthermore, by 8, $\mu(v_{b})$ does not belong to a pair of tree-clones. Finally, all leaves are bridge-nodes, therefore all unit vectors in $\mu(\mathcal{N})$ belong to bridge-nodes, so we only still have to consider $\mu$-vectors which are not unit vectors. Note that we also do not need to consider the unique maximal $\mu$-vector in $\mu(\mathcal{N})$.
Observation 5.
Now let $\mathcal{N}=(V,E)$ be a semi-binary stack-free network. Proposition 3 gives a set of conditions on the $\mu$-representation $\mu(\mathcal{N})$ which determine whether a given $\mu$-vector belongs to a bridge-node.
Proposition 3.
  1. 1.
  2. 2.
    one of the following holds
    1. (a)
      at least one of $\mu(k)$ and $\mu(\ell)$ has multiplicity 1 in $\mu(\mathcal{N})$; or
    2. (b)
      $\mu(k)=\mu(\ell)$ and #$\#\mu(k)=2$, and
  3. 3.
Proof.
For a more formal proof see the appendix. Here we would like to give a proof by illustration. First we will show that the three conditions hold for any bridge node $v_{b}$.
Figure 11: A semi-binary stack-free network with bridge-node $v_{b}$.
Observe Figure 11 for an example of condition 1. Note that none of the red ‘outside’ nodes have $\mu$-vectors which could ever contribute to a pair whose $\mu$-vectors sum up to $\mu(v_{b})$ because their $\mu$-vectors are simply too large and/or contain non-zero coordinates that correspond to leaves that are not below $v_{b}$. Note also that none of the $\mu$-vectors of the blue ‘inside’ nodes (which are not children of $v_{b}$) could ever contribute to a sum of only two $\mu$-vectors which sum up to $\mu(v_{b})$ because they are simply too small. Only the green child nodes of $v_{b}$ have $\mu$-vectors which together sum up to $\mu(v_{b})$. There is always just a single pair of $\mu$-vectors which sum up to $\mu(v_{b})$ if $v_{b}$ is a bridge-node and they belong to the children of $v_{b}$.
For the second condition consider what it means if either (a) or (b) would not hold. Then the children of $v_{b}$ have distinct $\mu$-vectors which both have multiplicity 2 or higher or they have equal $\mu$-vectors and the multiplicity is 3 or higher. There are three such possible cases: either both children are reticulations (Figure 12), or one of the children is a reticulation and the other is a tree-clone (Figure 13), or both are tree-clones (Figure 14). In the final case, they either have distinct $\mu$-vectors, or they have equal $\mu$-vectors and there is a third node with the same $\mu$-vector.
Figure 12: Two networks in which $v_{b}$ has two reticulation children. The dashed lines indicate that they connect somewhere above.
Figure 12 shows that even if one of the reticulation children of $v_{b}$ is a descendant of the other, there is a reticulation edge going into one of the reticulations (the red dashed line) which is not coming from $v_{b}$ and therefore $v_{b}$ is not a bridge-node.
Figure 13: Three networks in which $v_{b}$ has a reticulation and a tree-clone child.
Figure 13 illustrates the case when $v_{b}$ has a reticulation and a tree-clone as children. Note that even if either child is a descendant of the other child, then there is still a path (via the red dashed line) to some of the descendants of $v_{b}$ which does not pass through $v_{b}$. Therefore, $v_{b}$ cannot be a bridge-node.
Figure 14: Two networks in which $v_{b}$ has two tree-clone children.
Now assume $v_{b}$ has two children which are tree-clones which either have the same $\mu$-vector with multiplicity 3 or higher, or have distinct $\mu$-vectors. Figure 14 illustrates that in either of these cases there must be a path (via the red dashed line) to descendants of $v_{b}$ which do not pass through $v_{b}$, which implies $v_{b}$ is not a bridge-node.
For the third condition we refer back to Figure 11. Let us assume that $v_{b}$ is a bridge-node but $\mu(\mathcal{N})$ does contain vectors $\mu(x),\mu(y),\mu(z)$, such that $\mu(z)\nleq\mu(v_{b})$, $\mu(z)=\mu(x)+\mu(y)$ and $\mu(x)<\mu(v_{b})$. Note that $\mu(z)$ has to belong to a red ‘outside’ node and if it belongs to a reticulation then it also belongs to its tree-node child. Now imagine that the $\mu$-vector of such an outside tree-node is the sum of the $\mu$-vector of a blue ‘inside’ node and another $\mu$-vector. Even if there are multiple pairs (of possible children) whose $\mu$-vectors sum up to the $\mu$-vector of $z$ this would imply $z$ has a child which is not an ancestor of $v_{b}$. Yet, there are paths from this child to the leaves below $v_{b}$. This would mean there is a path to a leaf below $v_{b}$, which passes through $z$ and its child, which does not pass through $v_{b}$. This contradicts our assumption that $v_{b}$ is a bridge-node. Therefore if $v_{b}$ is a bridge-node, then condition 3 must hold.
To prove the other direction, we will show that if $v_{b}$ is not a bridge-node then one of the above conditions does not hold. If $v_{b}$ is not a bridge-node, then there is a path from the root to a descendant of $v_{b}$ which does not pass through $v_{b}$. Now because $\mu(\rho)\not<\mu(v_{b})$ and for any descendant $d$ of $v_{b}$, with $d\neq v_{b}$, we have $\mu(d)<\mu(v_{b})$. This implies that there is a node $z$ on that path with $\mu(z)\not<\mu(v_{b})$ which is the parent of a node $x$ with $\mu(x)<\mu(v_{b})$ (because $v_{b}$ is not a bridge-node this does not necessarily mean that $x$ is itself a descendant of $v_{b}$ but that does not matter). There are two cases to consider: either $\mu(z)=\mu(v_{b})$ or $\mu(z)\nleq\mu(v_{b})$.
Figure 15: Four networks in which $z$ and $v_{b}$ are tree-clones with $\mu(z)=\mu(v_{b})$.
If $\mu(z)=\mu(v_{b})$, then $\mu(v_{b})$ is a tree-clone. Figure 15 shows four networks where $v_{b}$ and $z$ share all, some, or none of their children. It can be seen that in the top network and the leftmost network where they share one or more of their children. Condition 2 of the proposition does not hold because the multiplicities of the $\mu$-vectors of the children are too high. In the third and fourth cases, they do not share any of their children. In the middle network, the children share the same pair of $\mu$-vectors and therefore the multiplicities are too high, again breaking condition 2. And in the rightmost network, there are two distinct pairs of $\mu$-vectors which sum up to the value of $\mu(v_{b})$ and $\mu(z)$, which means that the first condition does not hold.
Figure 16: Two networks where there exists a node $z$ with $\mu(z)\nleq\mu(v_{b})$ with a child $x$ such that $\mu(x)<\mu(v_{b})$.
Finally if $\mu(z)\nleq\mu(v_{b})$ then condition 3 does not hold because $\mu(z)=\mu(x)+\mu(y)$ for some node $x$ with $\mu(x)<\mu(v_{b})$. Figure 16 shows two such networks. The second network shows that $x$ does not have to be a descendant of $v_{b}$ in this case. We have now shown that if $v_{b}$ is not a bridge node then one of the conditions in Proposition 3 does not hold. We can conclude that the conditions in Proposition 3 are both necessary and sufficient for $\mu(v_{b})$ to belong to a bridge-node.
3 shows that for two semi-binary stack-free networks $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, if $\mu(\mathcal{N}_{1})=\mu(\mathcal{N}_{2})$, then $\mu(v_{b})\in\mu(\mathcal{N}_{1})$ belongs to a bridge-node in $\mathcal{N}_{1}$ if, and only if it belongs to a bridge-node in $\mathcal{N}_{2}$. In other words, the same $\mu$-vectors will belong to bridge-nodes in both networks. While $\mu(\rho)$ also satisfies the conditions of 3, a $\mu$-vector belonging to a bridge-node in $\mathcal{N}_{1}$ will never belong to the root in $\mathcal{N}_{2}$, since we can identify $\mu(\rho)$ as we noted in 5. 12 shows that if $\mu(v_{b})$ belongs to a bridge-node in both networks, then the same $\mu$-vector belongs to the lowest reticulation above that bridge-node in both networks. Then, by 4, these reticulations will be stable with respect to the same leaves. This leads us to the following theorem.
Theorem 3.
Proof.
As $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are strongly reticulation-visible, there is a tree-path to a bridge-node from the child of each reticulation. As mentioned above, the same $\mu$-vectors belong to bridge-nodes in both networks, and the same $\mu$-vectors belong to the lowest reticulation above those bridge-nodes. Therefore, the same $\mu$-vectors in both $\mu$-representations will belong to reticulations which are stable for the same set of leaves. To make this more clear, note the following. By 8, $\mu$-vectors which belong to stable nodes cannot belong to tree-clones. Furthermore, by 2, $\mu$-vectors with multiplicity greater than 2 do not belong to stable nodes. Moreover, as $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are strongly reticulation-visible, all the reticulations in both networks are stable. Therefore, the $\mu$-vectors with multiplicity greater than 2 in $\mu(\mathcal{N}_{1})$ and $\mu(\mathcal{N}_{2})$ do not belong to reticulations in either network. By combining 3 and 12, we can obtain the set of $\mu$-vectors with multiplicity 2 that is in bijection with the set of reticulations in both networks. In conclusion, the same $\mu$-vectors belong to reticulations in $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, and they are stable for a common set of leaves. Thus, by 2, $\bar{\mu}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{2})$. ∎
We now combine Theorem 1, 2, 3, and 12 to prove the following main result.
Theorem 4.
Proof.
Because $\mathcal{N}_{1}$ is orchard, it contains no tree-clones. Therefore, each $\mu$-vector in $\mu(\mathcal{N}_{1})$ has multiplicity at most 2. And the subset of $\mu(\mathcal{N}_{1})$ of $\mu$-vectors with multiplicity 2 is exactly the set of $\mu$-vectors which belong to reticulations in $\mathcal{N}_{1}$. Furthermore, because $\mathcal{N}_{1}$ is strongly reticulation-visible, these $\mu$-vectors belong to reticulations which are lowest above some bridge in $\mathcal{N}_{1}$. By 3, the same $\mu$-vectors belong to bridge-nodes in $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$. And by 12, every lowest reticulation - bridge-node pair is preserved in $\mathcal{N}_{2}$. As both $\mu$-representations do not contain vectors with multiplicity greater than 2, there are no other $\mu$-vectors belonging to reticulations in $\mathcal{N}_{2}$. Note that each $\mu$-vector with multiplicity 2 belongs to exactly one reticulation in $\mathcal{N}_{1}$. Therefore, $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ have the same reticulation set $R$ and each reticulation is stable for a common set of leaves in both networks. This means that, by 2, $\bar{\mu}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{2})$. Therefore, by Theorem 1, $\mathcal{N}_{1}\cong\mathcal{N}_{2}$. ∎
5 Conclusion and Discussion
Given class $\mathcal{N}_{1}$ class $\mathcal{N}_{2}$ Result Theorem
Table 1: A table detailing the main results. SB stands for semi-binary, SF stands for stack-free and SRV stands for strongly reticulation-visible. Note that strongly reticulation-visible networks are stack-free. The first row reads: “Given two networks $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ with the same modified $\mu$-representation, if $\mathcal{N}_{1}$ is semi-binary stack-free orchard and $\mathcal{N}_{2}$ is semi-binary stack-free, then they are isomorphic”.
We have shown that semi-binary stack-free orchard networks are encoded in the space of semi-binary stack-free networks by a modified $\mu$-representation (Theorem 1). This modified $\mu$-representation, contains the same path multiplicity vectors as the standard $\mu$-representation as originally proposed by Cardona et al. in [CRV08], but additionally includes the in-degrees of nodes. With this result we have shown that the cardinality of the symmetric difference of the $\bar{\mu}$-representations is a metric for the class of semi-binary stack-free orchard networks. We have also shown that this encoding result does not extend to non-binary stack-free orchard networks (Theorem 2) even if the outdegrees are also included in the modified $\mu$-representation.
We now give potential future research directions. All of the following points are elaborated on in [Rei23]. In [CPRC23] Cardona et al. propose an extended $\mu$-representation for encoding binary orchard networks. We wonder if this can be extended to encode semi-binary orchard networks. In another direction, we build on the results of Section 4.2. We showed that the bridge-nodes and stability of reticulations in semi-binary strongly reticulation-visible networks can be identified from their $\mu$-representations. Can we do the same for the non-binary variant? Finally, the original problem of encoding semi-binary stack-free orchard networks using $\mu$-representations remains open but perhaps our results here can illuminate the next steps towards a proof.
Appendix A Proof of Lemma 3
Proof.
First let us assume that $(b,a)$ is a cherry in $\mathcal{N}$. Let $p_{ab}$ be the parent of $a$ and $b$ in $\mathcal{N}$. In this case we have to show that $\bar{\mu}(\mathcal{N}^{\prime})$ contains exactly the $\bar{\mu}$-vectors $(\bar{\mu}(v)_{i})_{i\in(\{0\}\cup X\setminus\{b\})}$ for each $\bar{\mu}(v)\in\bar{\mu}(\mathcal{N})\setminus\{\bar{\mu}(b),\bar{\mu}(p_{ab})\}$, where $\bar{\mu}(p_{ab})=[1]\oplus(\mu(a)+\mu(b))$. First note that reducing $(b,a)$ in $\mathcal{N}$ does not change the number of paths to any leaf but $b$ for any node which is in both $\mathcal{N}$ and $\mathcal{N}^{\prime}$. Moreover, because $\mathcal{N}^{\prime}$ is a network on leaf set $X\setminus\{b\}$, for each node $v$ which is in both networks, $\bar{\mu}(\mathcal{N}^{\prime})$ contains $(\bar{\mu}(v)_{i})_{i\in\{0\}\cup X\setminus\{b\}}$. Furthermore, only the nodes $b$ and $p_{ab}$ are removed, when reducing $(b,a)$ in $\mathcal{N}$. Therefore $V^{\prime}=V\setminus\{b,p_{ab}\}$ and $p_{ab}$ is a tree-node with children $a$ and $b$ in $\mathcal{N}$, which means $\bar{\mu}(p_{ab})=\{1\}\oplus(\mu(a)+\mu(b))$ in $\bar{\mu}(\mathcal{N})$. This shows the claim and therefore $\bar{\mu}^{\prime}(\mathcal{N})=\bar{\mu}(\mathcal{N}^{\prime})$.
Now let us assume that $(b,a)$ is a simple reticulated cherry in $\mathcal{N}$, where $b$ has a reticulation parent $p_{b}$ and $p_{a}$ is the parent of $a$. To show that $\bar{\mu}^{\prime}(\mathcal{N})=\bar{\mu}(\mathcal{N}^{\prime})$, we have to show that $\bar{\mu}(\mathcal{N}^{\prime})$ contains the vector $\bar{\mu}(a)$ and for each $\bar{\mu}(v)$ in $\bar{\mu}(\mathcal{N})\setminus\{\bar{\mu}(p_{a}),\bar{\mu}(p_{b}),\bar{\mu}(a)\}$, $\bar{\mu}(\mathcal{N}^{\prime})$ contains a vector $\bar{\mu}^{\prime}(v)$ such that $\bar{\mu}^{\prime}(v)_{i}=\bar{\mu}(v)_{i}$ for $i\in X\setminus\{b\}$ and $\bar{\mu}^{\prime}(v)_{b}=\bar{\mu}(v)_{b}-\bar{\mu}(v)_{a}$. Note that in reducing $(b,a)$ in $\mathcal{N}$ the nodes $p_{a}$ and $p_{b}$ are suppressed therefore $V^{\prime}=V\setminus\{p_{a},p_{b}\}$. Furthermore, for each node in $V^{\prime}$ the number of paths to any leaf other than $b$ in $\mathcal{N}^{\prime}$ is equal to the number of paths to that leaf in $\mathcal{N}$. However, the paths to leaf $b$ which include $p_{a}$ in $\mathcal{N}$ are not present in $\mathcal{N}^{\prime}$ because of the removal of the edge $p_{a}p_{b}$. For any node other than $a$ the number of paths to leaf $b$ which pass through $p_{a}$ in $\mathcal{N}$ is equal to the number of paths to leaf $a$, but there are no paths from $a$ to $b$ in either network. Therefore, $\bar{\mu}(a)$ is an element of $\bar{\mu}(\mathcal{N}^{\prime})$ and for each other node $v$ which is in both networks $\bar{\mu}(\mathcal{N}^{\prime})$ contains the vector $\bar{\mu}^{\prime}(v)$ such that $\bar{\mu}^{\prime}(v)_{i}=\bar{\mu}(v)_{i}$ for $i\in X\setminus\{b\}$ and $\bar{\mu}^{\prime}(v)_{b}=\bar{\mu}(v)_{b}-\bar{\mu}(v)_{a}$. Thus proving the claim.
Finally, let us assume that $(b,a)$ is a complex reticulated cherry in $\mathcal{N}$. The only difference as compared to the case where $(b,a)$ was a simple reticulated cherry is the fact that when reducing $(b,a)$ in $\mathcal{N}$ the parent $p_{b}$ of $b$ is not suppressed but its in-degree is lowered by 1. Therefore, $\bar{\mu}(p_{b})$ is not removed from $\bar{\mu}(\mathcal{N})$ but $\bar{\mu}(p_{b})_{0}$ is lowered by 1. All other arguments still hold, thus also in this case $\bar{\mu}(\mathcal{N}^{\prime})=\bar{\mu}^{\prime}(\mathcal{N})$. ∎
Appendix B Formal proof of Proposition 3
We use the following lemmas (13 to 18) to show how to determine whether a non-unit $\mu$-vector belongs to a bridge-node (3).
@@ Appendix B Formal proof of Proposition 3
Given a tree-node $v_{b}$, let us first assume $v_{b}$ has two children $c_{1}$ and $c_{2}$ which are reticulations. Because $c_{1}$ and $c_{2}$ are reticulations they must have at least one other parent besides $v_{b}$. Neither $c_{1}$ is the parent of $c_{2}$ nor is $c_{2}$ the parent of $c_{1}$ because the network is stack-free. Note that, not both $c_{1}$ and $c_{2}$ can have a parent which is a descendant of the other one, because in that case there would be a cycle, from $c_{1}$ to the parent of $c_{2}$ to $c_{2}$ to the parent of $c_{1}$ to $c_{1}$. Furthermore, all descendants of $v_{b}$ except $v_{b}$ itself are descendants of $c_{1}$ or $c_{2}$. Therefore, either $c_{1}$ or $c_{2}$ has a parent which is not a descendant of $v_{b}$. But then there would be a path from the root to a leaf below $v_{b}$ via this parent, which does not pass through $v_{b}$. Thus, by 2, this means $v_{b}$ is not a bridge-node.
@@
We will show the lemma is true with a proof by contradiction, by showing that there cannot be a second pair. Assume $\mu(v_{b})$ is a bridge-node which is not a leaf and $\mu(\mathcal{N})$ contains at least two pairs $\mu(x),\mu(y)$ and $\mu(k),\mu(\ell)$ such that $\mu(v_{b})=\mu(x)+\mu(y)=\mu(k)+\mu(\ell)$. W.l.o.g. we can assume $\mu(x)$ and $\mu(y)$ belong to the children $x,y$ of $v_{b}$. By 13, we know that $\mu(k)\neq\mu(x)$, because otherwise $\mu(\ell)=\mu(y)$, in which case both #$\#\mu(x)\geq 2$ and #$\#\mu(y)\geq 2$, and if $\mu(x)=\mu(y)$ then #$\#\mu(x)\geq 4$. The same goes for $\mu(k)\neq\mu(y)$, $\mu(\ell)\neq\mu(x)$ and $\mu(\ell)\neq\mu(y)$. Note also that none of $x,y,k,\ell$ are ancestors of $v_{b}$ because their $\mu$-vectors are each lower than $\mu(v_{b})$. Now note that, by 3, $\mu(v_{b})>\mu(k)$ and $\mu(v_{b})>\mu(\ell)$ implies they are descendants of $v_{b}$. However, as we mentioned $x,y$ are the children of $v_{b}$. This means $k,\ell$ must be descendants of $x,y$, which are not equal to $x,y$, because their $\mu$-vectors differ. From this it follows that $\mu(x)+\mu(y)>\mu(k)+\mu(\ell)$. This contradicts our assumption that the sums were equal. ∎
@@
If $\mu(v_{t})\in\mu(\mathcal{N})$ belongs to a tree-clone, then there are at least two tree-nodes $v_{1}$ and $v_{2}$ with $\mu$-vector equal to $\mu(v_{1})$. By 1, unit-vectors do not belong to tree-clones, therefore $v_{1}$ and $v_{2}$ are not leaves. This means $v_{1}$ and $v_{2}$ each have two children. Let $c_{1}$ and $c_{2}$ be the children of $v_{1}$, and let $c_{3}$ and $c_{4}$ be the children of $v_{2}$. Note that $c_{1}$ must be distinct from $c_{2}$, and $c_{3}$ must be distinct from $c_{4}$, because phylogenetic networks do not contain parallel edges. We then have: $\mu(c_{1})+\mu(c_{2})=\mu(v_{1})=\mu(v_{2})=\mu(c_{3})+\mu(c_{4})$. The statement can be broken down into three unique cases.
@@
Let $\mu(v_{b})\in\mu(\mathcal{N})$ belong to a bridge-node $v_{b}$. Let $I_{b}$ be the set of leaves below $v_{b}$, so that $\mu(v_{b})_{i}=0$ for $i\notin I_{b}$. For any $\mu(x)\in\mu(\mathcal{N})$ with $\mu(x)\not<\mu(v_{b})$, it holds that $\mu(x)=P_{xv_{b}}\mu(v_{b})+\mu^{\prime}(x)$, where $P_{xv_{b}}$ is a non-negative integer, equal to the number of paths from $x$ to $v_{b}$, and $\mu^{\prime}(x)$ is a $\mu$-vector such that $\mu^{\prime}_{i}(x)=0$ for $i\in I_{b}$.
@@
Let $\mu(v_{b})\in\mu(\mathcal{N})$ belong to a bridge-node. For any $\mu(x)\in\mu(\mathcal{N})$ with $\mu(x)\not<\mu(v_{b})$, by 3, $\mu(x)$ does not belong to a descendant of the bridge-node with $\mu$-vector $\mu(v_{b})$. Therefore, all paths from $x$ to leaves below $v_{b}$ must visit $v_{b}$. Thus, if $a$ is a leaf below $v_{b}$, then any path from $x$ to $a$ is the composition of a path from $x$ to $v_{b}$ and a path from $v_{b}$ to $a$. It follows that for each such leaf $a$, $\mu(x)_{a}=P_{xv_{b}}\mu(v_{b})_{a}$. From this it follows that $\mu(x)=P_{xv_{b}}\mu(v_{b})+\mu^{\prime}(x)$, where $\mu^{\prime}(x)$ contains only the paths to leaves not below $v_{b}$, and therefore $\mu^{\prime}(x)$ is a $\mu$-vector such that $\mu^{\prime}_{i}(x)=0$ for $i\in I_{b}$. ∎
@@
We will show this with a proof by contradiction. Let us assume that $\mu(v_{b})\in\mu(\mathcal{N})$ belong to a bridge-node $v_{b}$ and $\mu(\mathcal{N})$ contains some $\mu(x),\mu(y)$ such that $\mu(x)<\mu(v_{b})$. Now let us assume, $\mu(\mathcal{N})$ contains $\mu(k),\mu(\ell)$ with $\mu(k)\not<\mu(v_{b})$ and $\mu(\ell)\not<\mu(v_{b})$, such that $\mu(k)+\mu(\ell)=\mu(x)+\mu(y)$. First note, that if $\mu(y)\leq\mu(v_{b})$, then $\mu(y)$ belongs to a descendant of $v_{b}$ and therefore $\mu(k)+\mu(\ell)=\mu(x)+\mu(y)\leq\mu(v_{b})$. But this contradicts our assumption that $\mu(k)\not<\mu(v_{b})$. Therefore, we can assume that $\mu(y)\nleq\mu(v_{b})$. Then, by 16:
@@
We will argue by contradiction. Assume $\mu(\mathcal{N})$ contains $\mu(v_{b}),\mu(z),\mu(x)$ and $\mu(y)$ as described in the lemma and let $\mu(v_{b})$ belong to a bridge-node $v_{b}$. If $\mu(z)$ belongs to a reticulation $r$, then the child of $r$ must be a tree-node with the same $\mu$-vector, so w.l.o.g. we can assume $\mu(z)$ belongs to a tree-node $z$. Because $\mu(z)=\mu(x)+\mu(y)$, we know $z$ is not a leaf, because $\mu(z)$ is not a unit vector. If $\mu(x),\mu(y)$ is the only pair in $\mu(\mathcal{N})$ which sum up to $\mu(z)$ then $z$ must be the parent of $x$. If there are more pairs in $\mu(\mathcal{N})$ then by 17, each of those pairs must contain at least one $\mu$-vector lower than $\mu(v_{b})$. This means that in any case $z$ will have a child with $\mu$-vector lower than $\mu(v_{b})$. However, $\mu(z)\nleq\mu(v_{b})$ implies that $z$ is not a descendant of $v_{b}$. Then there must be a path from the root to a leaf below $v_{b}$ via $z$ which does not visit $v_{b}$. By 2, this means that $v_{b}$ is not a bridge-node. Which contradicts our assumption. ∎
Proposition (3).
@@
  2. 2.
    one of the following holds
    1. (a)
      at least one of $\mu(k)$ and $\mu(\ell)$ has multiplicity 1 in $\mu(\mathcal{N})$; or
    2. (b)
      $\mu(k)=\mu(\ell)$ and #$\#\mu(k)=2$, and
  3. 3.